OpenStudy (anonymous):

PLEAAAASE HELP MEEEE Solve the following system of equations. -2x + 4y = 70 x + 5y = 42

OpenStudy (timmyk14):

Are x and y going to be the same value in these?

OpenStudy (anonymous):

@Timmyk14 what do you mean

OpenStudy (timmyk14):

Will the value of x in the equation, -2x + 4y = 70, will that value of x be the same value for the variable x in the equation x + 5y = 42

OpenStudy (anonymous):

@Timmyk14 yes i think so

OpenStudy (timmyk14):

ok, give me a minute to work this out (^_^)

OpenStudy (timmyk14):

x = -13 y = 11

OpenStudy (timmyk14):

-2(-13) = 26 26 + 4(11) = 70 -13 + 5(11) -13 + 55 +42

OpenStudy (timmyk14):

do you understand?

OpenStudy (timmyk14):



OpenStudy (anonymous):

@Timmyk14 thank you so much & yes i do understand

OpenStudy (timmyk14):

Ok (O_O)